OpenStudy (anonymous):

can someone help me with this problem Simplifying expressions with absolute values |-5| - 6|2| - |-1|

OpenStudy (anonymous):

it is 5-12-1 = 5-13 = -8

OpenStudy (anonymous):

did you understand how I arrived at the answer?

OpenStudy (anonymous):

i was jst about to ask you that.. but no i didnt.. can you explain

OpenStudy (anonymous):

what is the absolute value of -5?

OpenStudy (anonymous):

do you know what we mean by absolute value?

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

okay good. and the absolute value of -1?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so |-5| = 5 and |-1| =1 whats the absolute value of 2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

good, so |2| = 2

OpenStudy (anonymous):

so, |-5| - 6*|2| - |-1| = ??

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

great! you got your answer.

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

you are welcome.